Define the majority carrier concentration in an n-type Si semiconductor in terms of
the conduction band edge energy
1
Expert's answer
2020-10-05T10:53:24-0400

The majority carrier concentration in an n-type Si semiconductor in terms of

the conduction band edge energy is


"n_0=N_C\text{exp}\bigg(\frac{E_F-E_C}{kT}\bigg),\\\space\\
N_C=2\bigg(\frac{2M_C^*\pi k T}{h^2}\bigg)."

"N_C" is the effective density of states at the conduction band edge;

"M_C^*" is the effective mass of electron in the conduction band.